OpenStudy (anonymous):

Which expressions show how 6 • 45 can be rewritten using the distributive property? Choose all answers that are correct. A. 6 • 40 + 6 • 5 B. 6 • 40 − 6 • 5 C. 6 • 50 – 6 • 5 D. 6 • 50 + 5

OpenStudy (michele_laino):

hint: \[\Large 6 \cdot 45 = 6 \cdot \left( {40 + 5} \right)\]

OpenStudy (anonymous):

@Michele_Laino the answers is a and c

OpenStudy (michele_laino):

correct!